OpenStudy (anonymous):

Graph the integrals and use geometric formulas to evaluate the areas. (Do not integrate) *p.s. I will put the integral using equation editor in my reply section

OpenStudy (anonymous):

\[ \int\limits_{-6}^{3} (1/3x + 3) dx\] A=h/2 (b1 +b2)

OpenStudy (anonymous):

what is this?

OpenStudy (anonymous):

my professor wanted me to graph this integral, and find the area

OpenStudy (anonymous):

whoa, did u just do that?

OpenStudy (anonymous):

wolfie did it

OpenStudy (anonymous):

uhh okay, but how did u know how to do it?

OpenStudy (anonymous):

you have to

OpenStudy (anonymous):

i'm just asking cuz i have an exam this thursday, and i would have to know how to do this by hand, which i don't :(

OpenStudy (anonymous):

plot y=1/3x +3 do it using Draw button

OpenStudy (anonymous):

is it possible to do this manually?

OpenStudy (anonymous):

graph y=1/3 x +3 start at (0,3) go up by 1 , right by 3

OpenStudy (anonymous):

makes sense so far...